# library.bib

### v2

@@ 2.7.2 Temperedness and strong convergence
The following definition and proposition can be found in [MAgeeDelaSalle, Section 5]. Fix a generating set of size $r$ for the free group $F_{r}.$
@@
To prove that a function is tempered, we can use the following proposition (stated here for free groups only), from [MAgeeDelaSalle, Proposition 6.3].
@@
That this proposition holds for free groups is a result of Haagerup’s inequality [Haagerup, Lemma 1.5], which asserts that free groups with their standard generating sets have the rapid decay property.
@@ 2.8 Schreier graphs
Let $G$ be a finite group acting on a finite set $X$ and let $g_{1},\dots,g_{r}\in G.$ The Schreier graph $\mathrm{Sch}\left(S_{n}\curvearrowright X,g_{1},\dots,g_{r}\right)$ is the $2r$–regular graph on $|X|$ vertices, consisting of vertex set $X$ and, for every vertex $x\in X,$ for each $g_{i},$ we add an edge between $x$ and $g_{i}x.$ Note that we allow multiple edges and loops.
@@ 2.9 Word maps
$\mathbb{E}_{w}\left[\chi\right]\overset{\mathrm{def}}{=}\mathop{\mathbb{E}}_{g_{1},\dots,g_{m}\in G}\left[\chi\left(w\left(g_{1},\dots,g_{m}\right)\right)\right].$
@@
$\mathbb{E}_{w}\left[\chi\right]$ is equivalent to $\mathop{\mathbb{E}}_{\mathrm{Hom}\left(F_{r},S_{n}\right)}\left[\chi\left(\phi(w)\right)\right]$, where $\phi\in\mathrm{Hom}(F_{r},S_{n})$ is chosen uniformly randomly by choosing uniformly random i.i.d. $\sigma_{1},\dots,\sigma_{r}\in S_{n}$ and mapping each $x_{i}\mapsto\sigma_{i},$ as is described in §1.1.
@@
The results of this section are almost completely analogous to those of [MAgeeDelaSalle, Section 7]. Throughout, $C$ will denote a constant that does not depend on anything, but that may change from line to line. We require our bound on the expected irreducible stable character of a $w$–random permutation from Theorem 1.5, as well as our formulation of this expected character as a ratio of two polynomials in $\frac{1}{n},$ from Theorem 4.1.
Fix any integer $k>0$ and define
@@
Then $\Sigma_{n,k}\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k},$ so that
@@
The following theorem follows immediately by combining Theorem 1.5 and Theorem 4.1.
@@
    If $w$ is not the identity and $l(w)\leq q,$ then $g_{q,k}\phi_{w}$ is a polynomial of degree $\leq D_{q}=3kq+q^{2}.$
@@
    If $w$ is not a proper power then, for all $i<2k,$
@@
$g_{L,k}(x)=\prod_{c=1}^{L}(1-cx)^{L}\left[\prod_{j=1}^{2k}\left(1-(j-1)x\right)\right]^{L}$
@@
    $\left|\left(\frac{1}{g_{L,k}}\right)^{(i)}(t)\right|\leq(2i)!!2^{2i+1}\left(L^{3}+L(2k-1)^{2}\right)^{i}.$
@@
Fix any $0\leq t\leq\frac{1}{2}\frac{1}{L^{2}(L+1)+L(2k-1)(2k)}.$ That $g_{L,k}(t)\leq 1$ follows immediately from the fact that
$|t|<\min\Big{\{}\frac{1}{L},\frac{1}{2k-1}\Big{\}}.$
@@
$\geq$ $\exp\left(-2\left(\sum_{c=1}^{L}Lct+\sum_{j=1}^{2k}L(j-1)t\right)\right)$
$=$ $\exp\left(-2Lt\left(1+\dots+L+1+\dots+2k-1\right)\right)$
$=$ $\exp\left(-2Lt\left(\frac{L(L+1)}{2}+\frac{(2k-1)2k}{2}\right)\right)$
$=$ $\exp\left(-t\left(L^{2}(L+1)+L(2k-1)(2k)\right)\right)$
@@
$t\leq\frac{1}{2}\frac{1}{L^{2}(L+1)+L(2k-1)(2k)}<\frac{-\log\left(\frac{1}{2}\right)}{L^{2}(L+1)+L(2k-1)(2k)}.$
@@
$g_{L,k}(t)=\prod_{z=1}^{L^{2}+2kL}\left(1-b_{z}t\right)$
and then differentiate using the Leibniz rule. The $i^{\mathrm{th}}$ derivative is a sum of terms in which some terms are derived once (so are equal to $-b_{z}$) and some are not derived. Each factor that is not derived is of the form $1-b_{z}t,$ which belongs to $(0,1].$ So we can bound the derivative using only the terms that are derived:
@@
$\leq$ $\sum_{s=0}^{i}\sum_{\underset{\mathrm{distinct}}{u_{1},\dots,u_{s}}}\prod_{\alpha=1}^{s}b_{u_{\alpha}}$
$\leq$ $\sum_{s=0}^{i}\left(\sum_{z=1}^{L^{2}+2kL}b_{z}\right)^{s}$
$=$ $\sum_{s=0}^{i}\left(\sum_{c=1}^{L}Lc+\sum_{j=1}^{2k}L(j-1)\right)^{s}$
$\leq$ $\sum_{s=0}^{i}\left(L^{3}+L(2k-1)^{2}\right)^{s}$
$\leq$ $(i+1)\left(L^{3}+L(2k-1)^{2}\right)^{i}.$
To prove Part 3, we use that the $i^{\mathrm{th}}$ derivative of $\frac{1}{g_{L,k}}$ can be written as a product of $(2i)!!$! terms of the form
±$\frac{\pm g_{L,k}^{(\alpha_{1})}\dots g_{L,k}^{(\alpha_{i})}}{g_{L,k}^{i+1}},$
where $\alpha_{1},\dots,\alpha_{i}\in\mathbb{Z}_{\geq 0}$ and $\alpha_{1}+\dots+\alpha_{i}=i.$
By Part 2, each term is bounded by
$\frac{\prod_{j}\left(\alpha_{j}+1\right)\left(L^{3}+L(2k-1)^{2}\right)^{\alpha_{j}}}{g_{L,k}^{i+1}}$
@@
$2^{i+1}\prod_{j}\left(\alpha_{j}+1\right)\left(L^{3}+L(2k-1)^{2}\right)^{\alpha_{j}},$
which is then bounded by
$2^{2i+1}\left(L^{3}+L(2k-1)^{2}\right)^{i},$
@@
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|\phi_{x}^{(i)}(t)\right|}{i!}\leq h(i,q)||x||_{C^{*}\left(F_{r}\right)},$
@@
$\underset{(\dagger)}{\leq}$ $\frac{1}{n^{k}}\dim\left(\Sigma_{n,k}\right)||x||_{C^{*}\left(F_{r}\right)}$
@@
The inequality ()($\dagger$)( † ) follows since the map $w\mapsto\Sigma_{n,k}\left(w\left(\sigma_{1},\dots,\sigma_{r}\right)\right)$ is a unitary representation of $F_{r}$ for every $\left(\sigma_{1},\dots,\sigma_{r}\right)\in S_{n}^{r}.$ So we have
$\left|\left|\Sigma_{n,k}\left(x\left(\sigma_{1},\dots,\sigma_{r}\right)\right)\right|\right|\leq||x||_{C^{*}\left(F_{r}\right)}$
almost surely, from which we can bound
$\left|\mathbb{E}_{x}\left[\mathrm{Tr_{\Sigma_{n,k}}}\right]\right|\leq\dim\left(\Sigma_{n,k}\right)||x||_{C^{*}\left(F_{r}\right)}.$
By Lemma 2.16 (and using Stirling’s formula), for any integer $j,$ we can bound
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|P^{(j)}(t)\right|}{j!}\leq 2\left(\frac{CD_{q}^{4}}{j^{2}}\right)^{j}\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$ (9)
We have
$\frac{\phi_{x}^{(i)}}{i!}=\sum_{j=0}^{\min(i,D_{q})}\begin{pmatrix}i\\ j\end{pmatrix}\frac{j!}{i!}\frac{P^{(j)}}{j!}\left(\frac{1}{g_{q,k}}\right)^{(i-j)}.$
Note that $i\leq 2k\leq D_{q}=q^{2}+3kq,$ so that $\min(i,D_{q})=i.$ With the bound (9) and Lemma 3.2 Part 3, for $t\in\left[0,\frac{1}{2D_{q}^{2}}\right]$, each term (besides $j\neq i$ ) is smaller than $\frac{1}{i}2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i},$ so that
$\frac{\left|\phi_{x}^{(i)}(t)\right|}{i!}\leq 2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i}+\frac{i}{i}2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i},$
which proves the lemma. ∎
@@
We want to show that $\psi_{i}$ is tempered and satisfies the polynomial bound property in Proposition 2.20.
@@
For every integer $i$ with $0\leq i\leq k$, there is a polynomial $P$ of degree $14k+4i+$ 1 such that $\left|\psi_{i}(x)\right|\leq P_{n}(q)\|x\|_{C^{*}\left(F_{r}\right)}$ for every $q$ and for every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right].$
@@
$\left|\frac{\phi_{x}^{(k+i)}(0)}{(k+i)!}\right|\leq h(k+i,q)\|x\|_{C^{*}\left(F_{r}\right)}.$ (10)
@@
$\leq$ $C_{i}(1+qn)^{4k+4i+1}\mathbb{P}\left(\psi_{i}(g_{n})\neq 0\right)$
$=$ $C_{i}(1+qn)^{4k+4i+1}\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)$
$\leq$ $C_{i}C_{\mu}(1+qn)^{4k+4i+1}n^{5}\rho(\mu)^{n}.$
The first inequality follows from Lemma 3.4. Indeed, $g_{n}\in\mathbb{C}_{\leq qn}\left[F_{r}\right],$ so there is some constant $C_{i}$ for which $|\psi(g_{n})|\leq C_{i}\left(1+qn\right)^{4k+4i+1}.$ The final inequality follows from Proposition 2.18. It follows that
@@
For any $\epsilon>0,$ we have $\sup_{q\geq 1}h(2k,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}k^{12+\epsilon}\right)^{k}.$
@@
$4\left(CD_{q}^{4}\right)^{2k}\exp\left(\frac{-q}{\log\left(2+q\right)^{2}}\right)\leq 4\left(C\left(q^{2}+3kq\right)^{4}\right)^{2k}\exp\left(-q^{\frac{b-1}{b}}\right)$ (11)
@@
$\left(C^{2}k^{12+\epsilon}\right)^{k}$ (12)
for sufficiently large $b$ [^2] To see this, differentiate the RHS of (11) to see that the maximum is obtained around $q=Ck^{\frac{b}{b-1}}$and substitute this in to $h(2k,q)$.
@@
by substituting $q=a$ in to $h(2k,q)$ and ignoring the exponential term (since it is $\leq 1$). This is less than (12) (when $C$ is large enough) from which the lemma follows. ∎
@@
$\mathop{\left|\mathbb{E}\left[\mathrm{Tr}\left(\Sigma_{n,k}\left(x(\sigma_{1},\dots,\sigma_{r})\right)-\tau(x)\mathrm{Id}\right)\right]-\sum_{i=0}^{k-1}\frac{\psi_{i}(x)}{n^{i}}\right|}$ (13)
$\leq$ $\frac{\left(C^{2}k^{12+\epsilon}\right)^{k}}{n^{k}}w(q)\|x\|_{C^{*}\left(F_{r}\right)}.$
@@
Since $n\geq kq,$ then by Part 1 of Theorem 3.1, the LHS of (13) is equal to
@@
$\frac{n^{k}}{n^{2k}(2k)!}\left|\phi_{x}^{(2k)}\left(\frac{1}{n}\right)\right|.$
@@
$\frac{h(2k,i)}{n^{k}}\|x\|_{C^{*}\left(F_{r}\right)}.$ (14)
@@
$2\dim\left(\Sigma_{n,k}\right)\|x\|_{C^{*}\left(F_{r}\right)}+\sum_{i=0}^{k-1}\frac{\left|\psi_{i}(x)\right|}{n^{i}}$
@@
$n^{k}\left[h(0,q)+\sum_{i=0}^{k-1}\frac{1}{n^{k+i}}h(k+i,q)\right]\|x\|_{C^{*}\left(F_{r}\right)}.$
@@
$\sup_{q\geq 1}h(2k,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}k^{12+\epsilon}\right)^{k}$
@@
Fix $\alpha<\frac{1}{12},$ say $\alpha=\frac{1}{12}-\epsilon^{\prime},$ with $0<\epsilon^{\prime}\leq\frac{1}{12}.$ For each $n,$ and for any $k\leq n^{\alpha}$, let $\Pi_{n,k}$ be a random representation of $F_{r}$ given by
@@
$\left|\mathbb{E}\mathrm{Tr}\left(\Pi_{n,k}(x)\right)-T_{n}(x)\right|\leq\epsilon_{n}w(q)\|x\|_{C^{*}\left(F_{r}\right)},$
where, for some $\epsilon$ satisfying $0<\epsilon<\frac{12\epsilon^{\prime}}{\frac{1}{12}-\epsilon^{\prime}}$, we have
$\epsilon_{n}=\frac{\left(C^{2}k^{12+\epsilon}\right)^{k}}{n^{k}}\leq\left(\frac{C^{2}}{n^{1-(12+\epsilon)\alpha}}\right)^{k}$
@@
$T_{n}(x)=\dim\left(\Sigma_{n,k}\right)\tau(x)+\sum_{i=0}^{k-1}\frac{\psi_{i}(x)}{n^{i}}.$
By Lemma 3.5, $T_{n}$ is tempered, since it is the finite sum of tempered functions. Additionally, $T_{n}$ satisfies the polynomial bound
$\left|T_{n}(x)\right|\leq P_{n}(q)\|x\|_{C^{*}\left(F_{r}\right)}.$
@@
$\mathbb{P}\left[\left\|\Pi_{n,k}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right]\leq C(z,\delta)\left(\frac{C^{2}}{n^{1-(12+\epsilon)\alpha}}\right)^{k}.$

### v3

@@ 2.7.2 Temperedness and strong convergence
The following definition and proposition can be found in [MdlS24, Section 5]. Fix a generating set of size $r$ for the free group $F_{r}.$
@@
To prove that a function is tempered, we can use the following proposition (stated here for free groups only), from [MdlS24, Proposition 6.3].
@@
That this proposition holds for free groups is a result of Haagerup’s inequality [Haa78, Lemma 1.5], which asserts that free groups with their standard generating sets have the rapid decay property.
@@ 2.8 Schreier graphs
Let $G$ be a finite group acting on a finite set $X$ and let $g_{1},\dots,g_{r}\in G.$ The Schreier graph $\mathrm{Sch}\left(S_{n}\curvearrowright X,g_{1},\dots,g_{r}\right)$ is the 2r2r–regular graph on $|X|$ vertices, consisting of vertex set $X$ and, for every vertex $x\in X,$ for each $g_{i},$ we add an edge between $x$ and $g_{i}x.$ Note that we allow multiple edges and loops and that the graph is undirected.
By taking $g_{1},\dots,g_{r}\in G$ i.i.d. uniformly at random, one obtains a random 2r2r–regular graph on $|X|$ vertices.
@@ 2.9 Word maps
$\mathbb{E}_{w}\left[\chi\right]\overset{\mathrm{def}}{=}\mathop{\mathbb{E}}_{g_{1},\dots,g_{r}\in G}\left[\chi\left(w\left(g_{1},\dots,g_{r}\right)\right)\right].$
@@
$\mathbb{E}_{w}\left[\chi\right]$ is equivalent to $\mathop{\mathbb{E}}_{\hom\left(F_{r},S_{n}\right)}\left[\chi\left(\phi_{n}(w)\right)\right]$, where $\phi_{n}\in\hom(F_{r},S_{n})$ is chosen uniformly randomly by choosing uniformly random i.i.d. $\sigma_{1},\dots,\sigma_{r}\in S_{n}$ and mapping each $x_{i}\mapsto\sigma_{i},$ as is described in §1.2.
@@
The results of this section are somewhat analogous to those presented in [MdlS24, Section 7]. Throughout, $C$ will denote a constant that does not depend on anything, but that may change from line to line. We require our bound on the expected irreducible stable character of a ww–random permutation from Theorem 1.6. Fix any integer $K>0$ and define
@@
We view $\Sigma_{n,K}\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k},$ so that
@@
The following theorem follows immediately by combining Theorem 1.6 and Proposition 4.1 – recall that, for each $\lambda\vdash K$, we can write
where $\deg\left(P_{w,\lambda,q}\right)\leq 3Kq+Kq^{2}$ and that, as a rational function in $n$, $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O(1)$ for proper powers and $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{K}}\right)$ for non–powers.
@@
    If $w$ is not the identity and $l(w)\leq q,$ then $g_{q,K}\phi_{w}$ is a polynomial of degree $\leq D_{q}=3Kq+Kq^{2}+K.$
@@
    If $w$ is not the identity or a proper power then, for all $i<2K,$
@@
$g_{L,K}(x)=\prod_{c=1}^{KL}(1-cx)^{L}\left[\prod_{j=1}^{2K}\left(1-(j-1)x\right)\right]^{L}$
@@
    $\left|\left(\frac{1}{g_{L,K}}\right)^{(i)}(t)\right|\leq(2i)!!2^{i+1}\left(K^{2}L^{3}+L(2K-1)^{2}\right)^{i}.$
@@
Fix any $0\leq t\leq\frac{1}{2}\frac{1}{KL^{2}(KL+1)+L(2K-1)(2K)}.$ That $g_{L,K}(t)\leq 1$ follows immediately from the fact that
$|t|<\min\Big\{\frac{1}{KL},\frac{1}{2K-1}\Big\}.$
@@
$\geq$ $\exp\left(-2\left(\sum_{c=1}^{KL}Lct+\sum_{j=1}^{2K}L(j-1)t\right)\right)$
$=$ $\exp\left(-2Lt\left(1+\dots+KL+1+\dots+2K-1\right)\right)$
$=$ $\exp\left(-2Lt\left(\frac{KL(KL+1)}{2}+\frac{(2K-1)2K}{2}\right)\right)$
$=$ $\exp\left(-t\left(KL^{2}(KL+1)+L(2K-1)(2K)\right)\right)$
@@
$t\leq\frac{1}{2}\frac{1}{KL^{2}(KL+1)+L(2K-1)(2K)}<\frac{-\log\left(\frac{1}{2}\right)}{KL^{2}(KL+1)+L(2K-1)(2K)}.$
@@
$g_{L,K}(t)=\prod_{z=1}^{KL^{2}+2KL}\left(1-b_{z}t\right)$
and then differentiate using the Leibniz rule. The $i^{\mathrm{th}}$ derivative is a sum of terms in which $i$ terms are derived once (so are equal to $b_{z}$) and the others are not derived. Each factor that is not derived is of the form $1-b_{z}t,$ which belongs to $(0,1].$ So we can bound the derivative using only the terms that are derived:
@@
$\leq$ $\sum_{\underset{\mathrm{distinct}}{u_{1},\dots,u_{i}}}i!\prod_{\alpha=1}^{i}b_{u_{\alpha}}$
$\leq$ $\left(\sum_{z=1}^{KL^{2}+2KL}b_{z}\right)^{i}$
$=$ $\left(\sum_{c=1}^{KL}Lc+\sum_{j=1}^{2K}L(j-1)\right)^{i}$
$\leq$ $\left(K^{2}L^{3}+L(2K-1)^{2}\right)^{i}.$
To prove Part 3, we use that the $i^{\mathrm{th}}$ derivative of $\frac{1}{g_{L,K}}$ can be written as a product of $(2i)!!$ terms of the form
$\frac{\pm g_{L,K}^{(\alpha_{1})}\dots g_{L,K}^{(\alpha_{i})}}{g_{L,K}^{i+1}},$
where $\alpha_{1},\dots,\alpha_{i}\in\mathbb{Z}_{\geq 0}$ and $\alpha_{1}+\dots+\alpha_{i}=i.$ By Part 2, each term is bounded by
$\frac{\prod_{j}\left(K^{2}L^{3}+L(2K-1)^{2}\right)^{\alpha_{j}}}{g_{L,K}^{i+1}}$
@@
$2^{i+1}\prod_{j}\left(K^{2}L^{3}+L(2K-1)^{2}\right)^{\alpha_{j}}=2^{i+1}\left(K^{2}L^{3}+L(2K-1)^{2}\right)^{i},$
@@
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|\phi_{x}^{(i)}(t)\right|}{i!}\leq h(i,q)\left\|x\right\|_{C^{*}\left(F_{r}\right)},$
@@
$\underset{(\dagger)}{\leq}$ $\frac{1}{n^{K}}\dim\left(\Sigma_{n,K}\right)\left\|x\right\|_{C^{*}\left(F_{r}\right)}$
@@
The inequality $(\dagger)$ follows since the map $w\mapsto\Sigma_{n,K}\left(w\left(\sigma_{1},\dots,\sigma_{r}\right)\right)$ is a unitary representation of $F_{r}$ for every $\left(\sigma_{1},\dots,\sigma_{r}\right)\in S_{n}^{r}.$ So we have
$\left\|\Sigma_{n,K}\left(x\left(\sigma_{1},\dots,\sigma_{r}\right)\right)\right\|\leq\left\|x\right\|_{C^{*}\left(F_{r}\right)}$
almost surely, and since we can bound the trace of a matrix by its norm multiplied by its size, we can bound
$\left|\mathbb{E}_{x}\left[\mathrm{Tr_{\Sigma_{n,K}}}\right]\right|\leq\dim\left(\Sigma_{n,K}\right)\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$
By Lemma 2.16, for any integer $j,$ we can bound
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|P^{(j)}(t)\right|}{j!}\leq\frac{2^{2j+1}D_{q}^{4j}}{j!(2j-1)!!}\left\|x\right\|_{C^{*}\left(F_{r}\right)}=\frac{2^{3j+1}D_{q}^{4j}}{(2j)!}\left\|x\right\|_{C^{*}\left(F_{r}\right)},$
which, using Stirling’s formula, is bounded above by
$2\left(\frac{CD_{q}^{4}}{j^{2}}\right)^{j}\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$ (8)
We then have
$\frac{\phi_{x}^{(i)}}{i!}=\frac{1}{i!}\sum_{j=0}^{\min(i,D_{q})}\begin{pmatrix}i\\ j\end{pmatrix}P^{(j)}\left(\frac{1}{g_{q,K}}\right)^{(i-j)}=\sum_{j=0}^{\min(i,D_{q})}\frac{P^{(j)}}{j!}\frac{1}{(i-j)!}\left(\frac{1}{g_{q,K}}\right)^{(i-j)}.$
Note that $i\leq 2K\leq D_{q}=Kq^{2}+3Kq+K,$ so that $\min(i,D_{q})=i.$ By Lemma 3.2 Part 3 , for $t\in\left[0,\frac{1}{2D_{q}^{2}}\right]$,
$\frac{1}{(i-j)!}\left|\left(\frac{1}{g_{q,K}}\right)^{(i-j)}(t)\right|\leq 2^{2(i-j)+1}\left(K^{2}q^{3}+q(2K-1)^{2}\right)^{i-j}.$
Combining this with (8) we obtain that, in the range $t\in\left[0,\frac{1}{2D_{q}^{2}}\right],$
$\left|\frac{\phi_{x}^{(i)}}{i!}\right|\leq$ $4\left\|x\right\|_{C^{*}\left(F_{r}\right)}\sum_{j=0}^{i}\left(\frac{CD_{q}^{4}}{j^{2}}\right)^{j}4^{(i-j)}\left(K^{2}q^{3}+q(2K-1)^{2}\right)^{i-j}$
$\leq$ $4\left(CD_{q}^{2}\right)^{i}\left\|x\right\|_{C^{*}\left(F_{r}\right)}\sum_{j=0}^{i}\left(\frac{D_{q}^{2}}{j^{2}}\right)^{j},$
where the final inequality follows from the fact that $K^{2}q^{3}+q(2K-1)^{2}\leq D_{q}^{2}.$ Then observe that, for each $j$,
$\frac{\left(\frac{D_{q}^{2}}{j^{2}}\right)^{j}}{\left(\frac{D_{q}^{2}}{(j+1)^{2}}\right)^{j+1}}=\left(\frac{j+1}{j}\right)^{2j}\left(\frac{j+1}{D_{q}}\right)^{2}\leq e^{2},$
so that $\sum_{j=0}^{i}\left(\frac{D_{q}^{2}}{j^{2}}\right)^{j}\leq(1+e^{2})^{i}\left(\frac{D_{q}^{2}}{i^{2}}\right)^{i},$ which proves the lemma. ∎
@@
We want to show that $\psi_{i}$ is tempered for each $i<K$ and that it satisfies the polynomial bound property in Proposition 2.20.
@@
For every integer $i$ with $0\leq i\leq K$, there is a polynomial $P$ of degree $4K+4i+1$ such that $\left|\psi_{i}(x)\right|\leq P_{n}(q)\left\|x\right\|_{C^{*}\left(F_{r}\right)}$ for every $q$ and for every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right].$
@@
$\left|\frac{\phi_{x}^{(K+i)}(0)}{(K+i)!}\right|\leq h(K+i,q)\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$ (9)
@@
$\leq$ $C_{i}(1+q^{2}n^{2})^{4K+4i+1}\mathbb{P}\left(\psi_{i}(g_{n})\neq 0\right)$
$=$ $C_{i}(1+q^{2}n^{2})^{4K+4i+1}\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)$
$\leq$ $C_{i}C_{\mu}(1+q^{2}n^{2})^{4K+4i+1}n^{5}\rho(\mu)^{n}.$
The first inequality follows from Lemma 3.4. Indeed, $g_{n}\in\mathbb{C}_{\leq qn}\left[F_{r}\right],$ so there is some constant $C_{i}$ for which $|\psi(g_{n})|\leq C_{i}\left(1+q^{2}n^{2}\right)^{4K+4i+1}.$ The final inequality follows from Proposition 2.18. It follows that
@@
For any $\epsilon>0,$ we have $\sup_{q\geq 1}h(2K,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}K^{20+\epsilon}\right)^{K}.$
@@
$4\left(CD_{q}^{4}\right)^{2K}\exp\left(\frac{-q}{\log\left(2+q\right)^{2}}\right)\leq 4\left(C\left(Kq^{2}+3Kq+K\right)^{4}\right)^{2K}\exp\left(-q^{\frac{b-1}{b}}\right)$ (10)
@@
$\left(C^{2}K^{20+\epsilon}\right)^{K}$ (11)
for sufficiently large bb. [^3] To see this, differentiate the RHS of (10) to see that the maximum is obtained around $q=CK^{\frac{b}{b-1}}$and substitute this into $h(2K,q)$.
@@
by evaluating $h(2K,a)$ and ignoring the exponential term (since it is $\leq 1$). This is less than (11) (when $C$ is large enough) from which the lemma follows. ∎
We will also need the following observation to be used in Lemma 3.7, as well as our proof of Theorem 1.10 – the map
$u:x\in\mathbb{C}\left[F_{r}\right]\mapsto\tau(x)\frac{\phi_{e}^{(K)}(0)}{K!}\in\mathbb{C}$
is tempered. Indeed, if we write $\dim\Sigma_{n,K}=a_{0}+a_{1}n+\dots+a_{k}n^{K},$ one sees that $\frac{\phi_{e}^{(K)}(0)}{K!}=a_{0},$ a constant (in fact, this constant is always $(-1)^{K}$), and our observation follows from the fact that $x\mapsto\tau(x)$ is tempered. Moreover, this map obviously satisfies the polynomial bound $|u(x)|\leq\|x\|_{C^{*}\left(F_{r}\right)}.$
@@
$\left|\mathbb{E}\left[\mathrm{Tr}\big(\Sigma_{n,K}\left(x(\sigma_{1},\dots,\sigma_{r})\right)-\tau(x)\mathrm{Id}\big)\right]+\tau(x)\frac{\phi_{e}^{(K)}(0)}{K!}-\sum_{i=0}^{K-1}\frac{\psi_{i}(x)}{n^{i}}\right|$ (12)
is bounded by
$\frac{\left(C^{2}K^{20+\epsilon}\right)^{K}}{n^{K}}w(q)\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$
@@
We have
$\mathbb{E}\left[\mathrm{Tr}\left(\tau(x)\mathrm{Id}\right)\right]=\tau(x)\dim\Sigma_{n,K}=\tau(x)n^{K}\sum_{i=0}^{K-1}\frac{\phi_{e}^{(i)}(0)}{i!n^{i}}+\tau(x)\frac{\phi_{e}^{(K)}(0)}{K!}$
and, since $n\geq Kq,$ by Part 3 of Theorem 3.1, we have
$\sum_{i=0}^{K-1}\frac{\psi_{i}(x)}{n^{i}}=n^{K}\sum_{i=K}^{2K-1}\frac{\phi_{x}^{(i)}(0)}{i!n^{i}}=n^{K}\left[\sum_{w\neq e}x(w)\sum_{i=0}^{2K-1}\frac{\phi_{w}^{(i)}(0)}{i!n^{i}}+\tau(x)\sum_{i=K}^{2K-1}\frac{\phi_{e}^{(i)}(0)}{i!n^{i}}\right].$
Combining these observations with Part 1 of Theorem 3.1, we see that the LHS of (12) is equal to
@@
$\frac{n^{K}}{n^{2K}(2K)!}\left|\phi_{x}^{(2K)}\left(\frac{1}{n}\right)\right|.$
@@
$\frac{h(2K,q)}{n^{K}}\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$ (13)
@@
$\left(2\dim\left(\Sigma_{n,K}\right)+1\right)\left\|x\right\|_{C^{*}\left(F_{r}\right)}+\sum_{i=0}^{K-1}\frac{\left|\psi_{i}(x)\right|}{n^{i}}$
@@
$n^{K}\left[h(0,q)+\sum_{i=0}^{K-1}\frac{1}{n^{K+i}}h(K+i,q)\right]\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$
@@
$\sup_{q\geq 1}h(2K,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}K^{20+\epsilon}\right)^{K}$
@@
Fix $\alpha<\frac{1}{20},$ say $\alpha=\frac{1}{20}-\epsilon^{\prime},$ with $0<\epsilon^{\prime}\leq\frac{1}{20}.$ For each $n,$ and for any $K\leq n^{\alpha}$, let $\Pi_{n,K}$ be a random representation of $F_{r}$ given by
@@
$\left|\mathbb{E}\mathrm{Tr}\left(\Pi_{n,K}(x)\right)-T_{n}(x)\right|\leq\epsilon_{n}w(q)\left\|x\right\|_{C^{*}\left(F_{r}\right)},$
where, for some $\epsilon$ satisfying $0<\epsilon<\frac{20\epsilon^{\prime}}{\frac{1}{20}-\epsilon^{\prime}}$, we have
$\epsilon_{n}=\frac{\left(C^{2}K^{20+\epsilon}\right)^{K}}{n^{K}}\leq\left(\frac{C^{2}}{n^{1-(20+\epsilon)\alpha}}\right)^{K}$
@@
$T_{n}(x)=\dim\left(\Sigma_{n,K}\right)\tau(x)-\tau(x)\frac{\phi_{e}^{(K)}(0)}{K!}+\sum_{i=0}^{K-1}\frac{\psi_{i}(x)}{n^{i}}.$
By Lemma 3.5, $T_{n}$ is tempered, since it is a finite sum of tempered functions. Additionally, $T_{n}$ satisfies the polynomial bound
$\left|T_{n}(x)\right|\leq P_{n}(q)\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$
@@
$\mathbb{P}\left[\left\|\Pi_{n,K}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right]\leq C(z,\delta)\left(\frac{C^{2}}{n^{1-(20+\epsilon)\alpha}}\right)^{K}.$